OpenStudy (anonymous):

A bag has 12 brown marbles and 6 red marbles. Half of the red marbles are made of plastic. A marble is selected at random from the bag. What is the probability that it is a red, plastic marble? Write your answer as a fraction in simplest form.

OpenStudy (anonymous):

HELPPPPP

OpenStudy (uri):

Ha! Statistics,I'm pretty new to this sbuject :D @Feliixxia96 Which grade? :)

OpenStudy (anonymous):

8th

OpenStudy (uri):

I'll tag people who are great at this: @amistre64 @ganeshie8 @hartnn when they get online,you'll get help :)

OpenStudy (anonymous):

there are 6 red marbles

OpenStudy (anonymous):

half are made of plastic so there are 3 plastic red marbles

OpenStudy (anonymous):

ok

OpenStudy (uri):

Oh yeah @satellite73 i forgot to tag him,He's a genius :p

OpenStudy (anonymous):

there are \(12+6=18\) marbles all together

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the probability you get a red plastic marble is the ratio of the number of red plastic marbles, to the total number of marbles

OpenStudy (anonymous):

i.e. it is \(\frac{3}{18}=\frac{1}{6}\)

OpenStudy (anonymous):

so we would solve it?

OpenStudy (anonymous):

or is that the answer?

OpenStudy (anonymous):

yes, that is the answer

OpenStudy (anonymous):

ok, thank you :)

OpenStudy (anonymous):

it says its wrong

OpenStudy (anonymous):

??

OpenStudy (anonymous):

i am sticking with \(\frac{1}{6}\)

OpenStudy (anonymous):

where did you get the 1 from?

OpenStudy (anonymous):

3 red plastic marbles 18 marbles total ratio is \(\frac{3}{18}=\frac{1}{6}\)

OpenStudy (anonymous):

There are 12 normal marbles 3 plastic marbles and 3 other red marbles. So in total there is 18 marbles. Satellite is accurate. You can only have 3 plastic marbles. that is 3/18 = 1/6

OpenStudy (anonymous):

okay i get you

OpenStudy (anonymous):

thank you